OpenStudy (anonymous):

Perform the indicated operation. (36m2 - 60m + 25) ÷ (6m - 5) = _____

OpenStudy (anonymous):

6m + 5 6m - 5 R. 50 6m + 15 R. 100 6m - 5

OpenStudy (anonymous):

@zepdrix Do I solve this like I did the other equation?

OpenStudy (anonymous):

(6m-5)^2/(6m-5)=6m-5

zepdrix (zepdrix):

I think we would have to do `Polynomial Long Division` for this problem. Have you learned about that yet?

OpenStudy (anonymous):

Yes, but I'm not very good at it and probably don't remember all of it.

zepdrix (zepdrix):

Or if you're able to factor the top as marss has done, that works out nicely also.

OpenStudy (anonymous):

@marsss Why is (6m-5) squared?

OpenStudy (anonymous):

Is the numerator aka the top?

zepdrix (zepdrix):

yes that's what i meant by top, my bad :)

OpenStudy (anonymous):

So I would factor out each term seperately? For instance, I'd factor 36m^2. Then 60m. Then 25?

OpenStudy (anonymous):

(M+1)(M+36)

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Mmmmm I'm not sure how to explain factoring. \(\large 36m^2-60m+25 \qquad=\qquad 6^2m^2-30m-30m+5^2\) \(\large 6\cdot -5=-30\) So 6 and -5 are the numbers we want, \(\large (6m-5)(6m-5)\) I could explain it a little better if we did long division. But this method is a little faster I suppose.

OpenStudy (anonymous):

Would it be too much to ask you to explain it in long division?

zepdrix (zepdrix):

sure, np c:

OpenStudy (anonymous):

Thanks!

zepdrix (zepdrix):